### v1

@@
Self-driving cars increasingly rely on deep neural networks to achieve human-like driving [phan2023driveirl, marcu2023lingoqa]. However, the opacity of such black-box motion planners makes it challenging for the human behind the wheel to accurately anticipate when they will fail [nikolaidis2012human, major2020expect, paleja2021utility], with potentially catastrophic consequences [Titcomb2019, Fang2023, templeton2024waymo]. Here, we introduce concept-wrapper network (i.e., CW-Net), a method for explaining the behavior of black-box motion planners by grounding their reasoning in human-interpretable concepts. We deploy CW-Net on a real self-driving car and show that the resulting explanations refine the human driver’s mental model of the car, allowing them to better predict its behavior and adjust their own behavior accordingly. Unlike previous work using toy domains or simulations [kenny2023towards, kim2023and, atakishiyev2024explainable], our study presents the first real-world demonstration of how to build authentic autonomous vehicles (AVs) that give interpretable, causally faithful explanations for their decisions, without sacrificing performance. We anticipate our method could be applied to other safety-critical systems with a human in the loop, such as autonomous drones and robotic surgeons. Overall, our study suggests a pathway to explainability for autonomous agents as a whole, which can help make them more transparent, their deployment safer, and their usage more ethical.
@@
Self-driving cars, Mental model alignment, Interpretable ML, AI safety
@@ 1 Introduction
There are hundreds of companies developing autonomous vehicle (AV) technology globally [badue2021self], promising to revolutionize transportation for everyone. However, the complexity of fully driverless autonomy has prompted an industry shift towards advanced driver-assistance systems, which require successful communication between the AV and human driver. This is made increasingly difficult by the adoption of deep neural networks in AVs for planning and decision making, the core cognitive functions that determine driving behavior  [atakishiyev2024explainable]. Deep learning allows motion planners to learn the nuances of human driving behavior from data, but the implicit nature of the learned driving policies makes it challenging to understand the causes of their decisions and to predict their behavior.
@@
A lack of effective communication between the AV and the human driver has contributed to multiple high-profile incidents, some resulting in fatalities [Titcomb2019, Fang2023, templeton2024waymo], highlighting the urgent need to make deep motion planners interpretable [atakishiyev2024explainable]. Previous studies have sought to address this need using surveys and simulated scenarios [koo2015did, koo2016understanding, wiegand2020d, wang2021human, schneider2021increasing, schneider2021explain, omeiza2021towards, zemni2023octet], a human driver emulating the AV [kim2023and, schneider2023don], or large language models providing post-hoc explanations [marcu2023lingoqa]. However, these studies were theoretical, did not provide faithful explanations of the AV’s reasoning process, or were only evaluated in simulation. This leaves open the question of how to provide understandable and useful explanations for the decisions of deep motion planners deployed in real self-driving cars.
@@
To answer this question, we scale up our recent work on interpretable-by-design deep reinforcement learning with prototype-wrapper networks (PW-Nets) [kenny2023towards], using motifs from the literature on concept-bottleneck models [yuksekgonulpost]. Our key proposal is to ground the reasoning of black-box motion planners in human-interpretable concepts, such as “Approaching stopped vehicle” or “Close to cyclist”. This method is rooted in case-based reasoning, a classical artificial intelligence (AI) approach [leake1996case, keane2019case, sormo2005explanation] inspired by cognitive models of human reasoning and memory [schank1983dynamic]. It results in causal explanations, such as “I chose to stop based on recognizing that we are approaching a stopped vehicle.”. In our tests, these explanations help align a safety driver’s mental model of the AV with its actual internal decision-making process, increasing transparency and predictability. Importantly, this approach can be applied to arbitrary pre-trained deep neural networks, does not require retraining from scratch, and does not degrade performance of the original black-box planner.
@@
We apply our proposed method, CW-Net (short for concept-wrapper network), to a deep motion planner trained to imitate human driving behavior using inverse reinforcement learning [phan2023driveirl]. We replace the final (reward) layer of the pretrained deep neural network with a concept classifier, followed by a new reward layer. We then jointly train the classifier and the new reward layer to predict scenario types and driving decisions, respectively, without modifying the rest of the network. Evaluation on a large-scale benchmark [karnchanachari2024towards] confirms that CW-Net is able to classify concepts without compromising driving behavior. To study the utility of the explanations, we then deploy CW-Net on a real self-driving car [heim2024lab2car]. We demonstrate three situations in which the driver has an inaccurate mental model of the motion planner, which is subsequently corrected by the explanations. This ultimately changes the behavior of the driver, for example, by increasing their vigilance in certain situations. Finally, we confirm the statistical significance of these results in an online study (N=120). Overall, our work demonstrates how explainable AI can help users of advanced autonomous systems better understand their behavior in naturalistic settings [endsley2017autonomous], while also providing insights that can potentially accelerate the development and refinement of such systems.
2 Black-box motion planner
@@
Figure 1: Planner architecture. a. Autonomous vehicle stack. Sensory input is processed by the perception module to generate scene context $s$. The planning module processes $s$ to compute trajectory $\hat{\tau}$, which is followed by the control module. b. Black-box motion planner. $s$ is fed to trajectory generator $G$, which produces candidate trajectories $\{\tau_{1}\dots\tau_{k}\}$, and scene encoder $H$, which produces a scene embedding $\mathbf{h}$. These are fed into encoder $E$, which produces scene-trajectory embeddings $\mathbf{z}_{i}$ for each $\tau_{i}$, which are in turn fed into the reward layer $R$. $R$ computes a reward $r_{i}$ for each trajectory. The model is trained to output higher rewards for trajectories closer to the the ground-truth human trajectories. c. CW-Net. Identical to b., except $\mathbf{z}_{i}$ is fed to a concept classifier $C$, which produces concept assignments $\mathbf{c}_{i}$. These are fed to a new reward layer $R^{\prime}$ to produce rewards $r^{\prime}_{i}$. In parallel, $\mathbf{c}_{i}$ is processed to generate the explanations. The reward layer is trained to prefer the same trajectories as the black-box planner, while the concept layer is supervised with ground-truth scenario labels. The weights of the faded components are frozen during training. CLOSE, “Close to another vehicle”. ASV, “Approaching stopped vehicle”. BIKE, “Close to cyclist”.
@@
We focus on the motion planning module of the AV stack (Figure 1a), which takes as input a scene context $s$ and outputs a trajectory $\hat{\tau}$. $s$ is a symbolic object-oriented representation of the scene computed by the perception module, while $\hat{\tau}$ is the trajectory that the subsequent controller module should follow. We use a deep neural network architecture consisting of a scene encoder $H(s)\rightarrow\mathbf{h}$ and a trajectory generator $G(s)\rightarrow\{\tau_{1}\dots\tau_{k}\}$, followed by a scene-trajectory encoder $E(\mathbf{h},\tau_{i})\rightarrow\mathbf{z}_{i}$ and a final reward layer $R(\mathbf{z}_{i})\rightarrow r_{i}$ (Figure 1b). $H$ computes a scene embedding $\mathbf{h}$ and $G$ computes a set of $k$ candidate trajectories $\{\tau_{1}\dots\tau_{k}\}$. Those are combined in $E$ to compute an embedding $\mathbf{z}_{i}$ for each trajectory $\tau_{i}$. Finally, $R$ computes an estimated reward $r_{i}$ for each trajectory $\tau_{i}$, quantifying how human-like it is. In other words, $r_{i}$ is higher when $\tau_{i}$ is more similar to how a human would drive in this situation.
@@
$\hat{\tau}=\tau_{\hat{i}}\quad\text{such that}\quad\hat{i}=\underset{i\in{1,...,k}}{\arg\max}\,r_{i},\quad r_{i}=R(E(\mathbf{h},\tau_{i}))$
We train the planner on 80 hours of human expert driving using inverse reinforcement learning [phan2023driveirl].
@@ 3 Planning over human-friendly concepts
To make the planner more interpretable, we replace $R$ with a concept classifier $C(\mathbf{z}_{i})\rightarrow\mathbf{c}_{i}$, followed by a new reward layer $R^{\prime}(\mathbf{c}_{i})\rightarrow r^{\prime}_{i}$ (Figure 1c). The concept classifier $C$ computes a logit vector $\mathbf{c}_{i}$ which is passed through a softmax and/or sigmoid layer which assigns probabilities to different human-interpretable concepts. Since $R^{\prime}$ computes trajectory rewards from $\mathbf{c}_{i}$, the final decisions are based solely on these concept assignments and hence they constitute a causally faithful explanation. The rest of the network remains the same.
@@
$\hat{\tau}=\tau_{\hat{i}}\quad\text{such that}\quad\hat{i}=\underset{i\in{1,...,k}}{\arg\max}\,r^{\prime}_{i},\quad r^{\prime}_{i}=R^{\prime}(C(E(\mathbf{h},\tau_{i})))$
@@
We train CW-Net to jointly predict concept labels and mimic the driving decisions of the black-box planner. Specifically, $\mathbf{c}_{i}$ is supervised with multinomial labels corresponding to types of scenarios, such as “Approaching stopped vehicle” or “Close to cyclist”. This ensures that CW-Net assigns a unique interpretable concept to each unit in $\mathbf{c}_{i}$. At the same time, $r^{\prime}_{i}$ is supervised with trajectories selected by the black-box planner using a cross-entropy loss. During training, the rest of the deep neural network ($H$, $G$, and $E$) is kept frozen (see Section 8.2 for more details).
4 Evaluation in simulation
@@
As a baseline, we first evaluated the black-box planner (without CW-Net) using closed-loop simulations on the nuPlan dataset [karnchanachari2024towards] (Table S1). Overall, the results were competitive with the top submissions to the nuPlan challenge, although performance was slightly lacking when starting from a stop (see Section Simulation Results). This suggests that there is room for improvement and, importantly, opportunities to study explanations of undesirable behavior.
@@
We then evaluated the driving performance of CW-Net wrapped around the black-box planner (Table S1). The results were equivalent, with less than 1% difference across all metrics, confirming our method did not degrade driving performance. We also evaluated concept classification on held-out datasets (Table S2/S3). Mean accuracy was 54%, with 23% precision, 77% recall, and an F1 score of 0.31 (see Section 8.3). Overall, these results indicate that CW-Net can be used to ground the decision making of high-performance deep motion planners in human-interpretable concepts without sacrificing driving performance.
5 Explanations in real-world deployment
@@
Figure 2: Deployment setup. A safety driver, a support engineer, and a researcher were present. The safety driver drove the AV manually between road tests, engaged self-driving mode at the start of each test, monitored AV performance during the test, and took over in case of unsafe driving. The support engineer deployed CW-Net and set scenario destinations. The researcher directed testing. The dashboard included a map with overlaid object detections ($s$) from the perception module and the output trajectory ($\hat{\tau}$). Explanations $\mathbf{c}_{\hat{i}}$ from CW-Net were shown as percentages for easier interpretation [gigerenzer1995improve].
To evaluate the usefulness of the explanations in naturalistic settings [endsley2017autonomous], we deployed CW-Net on a real AV using the Lab2Car wrapper [heim2024lab2car]. All tests were performed on a closed course or private lot with an experienced safety driver. Figure 2 shows the experimental setup inside the AV. We next detail three notable situations in which the explanations proved beneficial.
@@ 5.1 Unexpected stopping for nearby vehicles
Figure 3: Results: a. The CLOSE concept activated when the car got stuck next to parked vehicles. The driver initially thought the pick-up/drop-off area was the cause, but the explanation suggested that it was the nearby vehicles. When the driver engaged self-driving away from the park vehicles, activation of the CLOSE concept decreased and the AV started moving again, counter to driver’s initial mental model and consistent with the explanations. Across tests, CLOSE correlated with speed and the intercept accurately predicted this event. b. The ASV concept activated when the car stopped next to a traffic cone. The driver initially thought the cone was the cause, but the explanation suggested that the AV was hallucinating a stopped vehicle. When we removed the cone in a counterfactual test, the same phantom braking and concept activation occurred. Across tests, ASV correlated with reductions in speed when spiking above 0.5 probability. c. The BIKE concept failed to activate in our first round of tests with a cyclist, but the car always stopped safely for the cyclist. Responding to the explanation, the driver engaged self-driving from slower speeds in a second round of tests. Follow-up analyses revealed that the AV stopped for the cyclist due to backup safety mechanisms unrelated to CW-Net, which indicates that the driver’s increased level of caution was appropriate. These three scenarios validate the driver’s updated mental models in response to the explanations. a-c. Findings were confirmed in an online study which replicated the events from the road. After seeing the explanations, subjects showed less agreement with the driver’s initial mental model, compared to a control group. Standard error of the mean and 3-second rolling averages is shown in relevant plots.
We observed that the AV repeatedly came to a stop shortly before a pedestrian pickup/drop-off zone (Figure 3a). The driver’s intuition was that the car stopped because of the pickup/drop-off zone, but the explanations indicated that the planner stopped because it detected that it was “Close to another vehicle” (the CLOSE concept). To test this hypothesis, the driver manually moved the car farther from the parked cars. At this point, the probability of CLOSE decreased and the AV began moving again, thus confirming the alternative hypothesis. A full timeline of events is detailed in Figure 3a. We fitted the intercept of the CLOSE probability against the speed of the AV globally and found it accurately predicts stopping and starting for this event. [^1] This situation used a variant of our architecture detailed in Figure S1.

### v2

@@
Self-driving cars increasingly rely on deep neural networks to achieve human-like driving [1, 2, 3]. The opacity of such black-box planners makes it challenging for the human behind the wheel to accurately anticipate when they will fail [4, 5, 6], with potentially catastrophic consequences [7, 8, 9]. While research into interpreting these systems has surged, most of it is confined to simulations or toy setups due to the difficulty of real-world deployment [10, 11], leaving the practical utility of such techniques unknown. Here, we introduce the Concept-Wrapper Network (CW-Net), a method for explaining the behavior of machine-learning-based planners by grounding their reasoning in human-interpretable concepts. We deploy CW-Net on a real self-driving car and show that the resulting explanations improve the human driver’s mental model of the car, allowing them to better predict its behavior. To our knowledge, this is the first demonstration that explainable deep learning integrated into self-driving cars can be both understandable and useful in a realistic deployment setting. CW-Net accomplishes this level of intelligibility while providing explanations which are causally faithful and do not sacrifice driving performance. Overall, our study establishes a general pathway to interpretability for autonomous agents by way of concept-based explanations, which could help make them more transparent and safe.
@@
Self-driving cars, Mental models, Situational awareness, Interpretable ML
@@ 1 Introduction
There are hundreds of companies developing autonomous vehicle (AV) technology globally [12], promising to revolutionize transportation for everyone. The industry currently spans two segments: driverless ride-hail systems and consumer vehicles [13]. In consumer vehicles, machine learning (ML) solutions have markedly improved advanced driver-assistance systems, yet they still require human intervention in unusual or challenging situations where learned planners may not determine the correct action [14]. As driving is safety-critical, such infrequent failures matter, making it essential that the human driver is able to anticipate and be prepared for such situations [15]. However, the opaque nature of ML planners makes it challenging to interpret and communicate the causes of their decisions, hampering the ability of human drivers to understand and predict AV behavior while achieving real-time situational awareness [16, 17, 18].
@@
Lack of effective communication between the AV and the human driver has contributed to multiple high-profile incidents, some resulting in fatalities [7, 8, 9], highlighting the urgent need to make ML planners interpretable [11]. Previous studies have sought to address this using surveys and simulated scenarios [19, 20, 21, 22, 23, 24, 25, 26], a human driver emulating the AV [27, 28], or large language models providing post-hoc explanations [3]. However, these studies were theoretical, did not provide causally faithful explanations, were only evaluated in simulation, or did not convincingly show the practical utility of the explanations. This leaves open the question of how to provide explanations that are understandable, useful, and faithful to the decision-making process of the AV in a realistic setting.
@@
To answer this question, we scale up our recent work on interpretable-by-design deep reinforcement learning [10] using motifs from the literature on concept-bottleneck models [29] to propose the Concept-Wrapper Network (CW-Net). CW-Net grounds the reasoning of a black-box ML planner in human-interpretable concepts, such as “Approaching stopped vehicle” or “Close to cyclist”. This method is rooted in case-based reasoning, a classical artificial intelligence (AI) approach [30, 31, 32] inspired by cognitive models of human reasoning and memory [33]. CW-Net can be applied to arbitrary pretrained deep neural networks, does not require retraining from scratch, and does not degrade the performance of the original black-box ML planner. Since the inferred concepts are the sole input to the final decision-making module of CW-Net that directly dictates AV behavior, we refer to them as causally faithful explanations.
@@
We apply CW-Net to a ML planner trained to imitate human driving behavior using inverse reinforcement learning [1, 2]. We replace the final (reward) layer of the pretrained deep neural network with a concept classifier, followed by a new reward layer. We then jointly train the classifier and the new reward layer to predict scenario types and driving decisions, respectively, without modifying the rest of the network. Evaluation on a large-scale benchmark [10] confirms that CW-Net is able to classify concepts without compromising driving behavior. To study the utility of the explanations, we deploy CW-Net on a real self-driving car with a safety driver in a semi-naturalistic study [35]. We demonstrate three situations in which the driver’s initially inaccurate mental model is subsequently improved by the CW-Net explanations, which in turn allows the driver to more accurately predict AV behavior. We then observe similar mental model improvement in a larger online study simulating these scenarios. Lastly, we link these results to more complex real-world scenarios by deploying CW-Net on public roads in Las Vegas. We use these new data in a large online study ($N=100$) to demonstrate that improvements in mental model goodness lead to improved predictive ability and situational awareness. Overall, our work demonstrates how explainable AI can help users of an advanced autonomous system improve their mental model of the system and better anticipate its future behavior.
2 Black-box planner
@@
Figure 1: Planner architecture, adapted with permission from Tomov et al. (2025) [1]. a. Autonomous vehicle stack. Sensory input is processed by the perception module to generate scene context $s$. The planning module processes $s$ to compute trajectory $\hat{\tau}$, which is followed by the control module. b. Black-box ML planner. Here, $s$ is fed to trajectory generator $G$, which produces candidate trajectories $\{\tau_{1}\dots\tau_{k}\}$, and scene encoder $H$, which produces a scene embedding $\mathbf{h}$. These are fed into encoder $E$, which produces scene-trajectory embeddings $\mathbf{z}_{i}$ for each $\tau_{i}$, which are in turn fed into the reward layer $R$. $R$ computes a reward $r_{i}$ for each trajectory. The model is trained to output higher rewards for trajectories closer to the the ground-truth human trajectories. c. CW-Net. Identical to b., except $\mathbf{z}_{i}$ is fed to a concept classifier $C$, which produces concept assignments $\mathbf{c}_{i}$. These are fed to a new reward layer $R^{\prime}$ to produce rewards $r^{\prime}_{i}$. In parallel, $\mathbf{c}_{i}$ is processed to generate the explanations. The reward layer is trained to prefer the same trajectories as the black-box ML planner, while the concept layer is supervised with ground-truth scenario labels. The weights of the faded components are frozen during training. CLOSE, “Close to another vehicle”. ASV, “Approaching stopped vehicle”. BIKE, “Close to cyclist”.
@@
We focus on the planning module of the AV stack (Figure 1a), which takes as input a scene context $s$ and outputs a trajectory $\hat{\tau}$. Here, $s$ is a symbolic object-oriented representation of the scene, computed by the perception module, while $\hat{\tau}$ is the trajectory that the subsequent controller module should follow. We use a deep neural network architecture [1, 2] consisting of a scene encoder $H(s)\rightarrow\mathbf{h}$ and a trajectory generator $G(s)\rightarrow\{\tau_{1}\dots\tau_{k}\}$, followed by a scene-trajectory encoder $E(\mathbf{h},\tau_{i})\rightarrow\mathbf{z}_{i}$ and a final reward layer $R(\mathbf{z}_{i})\rightarrow r_{i}$ (Figure 1b). $H$ computes a scene embedding $\mathbf{h}$ and $G$ computes a set of $k$ candidate trajectories $\{\tau_{1}\dots\tau_{k}\}$. Those are combined in $E$ to compute an embedding $\mathbf{z}_{i}$ for each trajectory $\tau_{i}$. Finally, $R$ computes an estimated reward $r_{i}$ for each trajectory $\tau_{i}$, quantifying how human-like it is. In other words, $r_{i}$ is higher when $\tau_{i}$ is more similar to how a human would drive in this situation.
@@
$\hat{\tau}=\underset{\tau\in{\{\tau_{1}...\tau_{k}}\}}{\arg\max}R(E(\mathbf{h},\tau))$
We use inverse reinforcement learning to train the planner on 80 hours of human expert driving [1, 2].
@@ 3 Planning over human-friendly concepts
To make the ML planner interpretable, we modify the planner architecture to additionally provide explanations for its behavior [10]. Our working definition of explainable AI follows Gunning & Aha [36], who define it as ‘AI systems that can explain their rationale to a human user, characterize their strengths and weaknesses, and convey an understanding of how they will behave in the future.’
Specifically, we replace $R$ with a concept classifier $C(\mathbf{z}_{i})\rightarrow\mathbf{c}_{i}$, followed by a new reward layer $R^{\prime}(\mathbf{c}_{i})\rightarrow r^{\prime}_{i}$ (Figure 1c). The concept classifier $C$ computes a logit vector $\mathbf{c}_{i}$ that is passed through a softmax and/or sigmoid layer that, in turn, assigns probabilities to different human-interpretable concepts. Since $R^{\prime}$ computes trajectory rewards from $\mathbf{c}_{i}$, the final decisions are based solely on these concept assignments and hence they constitute a causally faithful explanation. The rest of the network remains the same.
@@
$\hat{\tau}=\underset{\tau\in{\{\tau_{1}...\tau_{k}}\}}{\arg\max}R^{\prime}(C(E(\mathbf{h},\tau)))$
@@
We train CW-Net to jointly predict concept labels and mimic the driving decisions of the black-box planner. Specifically, $\mathbf{c}_{i}$ is supervised with labels corresponding to types of scenarios, such as “Approaching stopped vehicle” or “Close to cyclist” (see Section 8.5 for a full list of concepts). This ensures that CW-Net assigns a unique interpretable concept to each unit in $\mathbf{c}_{i}$. At the same time, $r^{\prime}_{i}$ is supervised with trajectories selected by the black-box planner using a cross-entropy loss. During training, the rest of the deep neural network ($H$, $G$, and $E$) is kept frozen (see Section 8.2 for more details). We focus our study on this particular ML planner architecture, noting that prior work indicates CW-Net would generalize effectively to other architectures [10].
4 Classifying concepts while preserving planner performance
@@
As a baseline, we first evaluated the black-box planner (without CW-Net) using closed-loop simulations with the nuPlan simulator on the nuPlan dataset [10] (Table S1). Overall, the results were competitive with the top submissions to the nuPlan challenge, although performance was slightly lacking when starting from a stop (see Section 8). This suggests that there is room for improvement and, importantly, opportunities to study explanations of undesirable behavior.
@@
We then evaluated the driving performance of CW-Net wrapped around the black-box planner (Table S1). The results were equivalent, with less than 1% difference across all metrics, confirming that our method did not degrade driving performance. We also evaluated concept classification on held-out datasets (Tables S2S3). Mean accuracy was 54%, with 23% precision, 77% recall, and an F1 score of 0.31 (see Section 8.3). Overall, these results indicate that CW-Net can be used to ground the decision making of high-performance ML planners in human-interpretable concepts, without sacrificing driving performance.
5 Mental model improvement in surprising real-world situations
Our central hypothesis is that the explanations from CW-Net would improve the human driver’s mental model of the AV and, by extension, their situational awareness. This would be particularly salient in surprising situations, which is when explanations are most useful [37]. Specifically, the explanations should improve the driver’s ability to understand and address the reasons for AV failures, and also to predict its future behavior.
To test this hypothesis, we deploy CW-Net on a real AV on a private track using the Lab2Car wrapper [35] and observe how safety drivers react to surprising events and the corresponding CW-Net explanations (Figure 2). These situations were not planned, but instead occurred naturally, with minimal intervention from the researchers, who only dictated high-level plans for each day. This semi-naturalistic study allowed us to assess the utility of the explanations in naturally occurring surprising situations. To measure AV predictability, we record the drivers’ ability to make counterfactual predictions about AV behavior in these situations. To note the drivers’ mental models, together with their predictive ability, we additionally record their think-aloud thoughts before and after considering CW-Net explanations [11]. We focus on concepts that relate to other road users and can be easily tested counterfactually (see Section 8.5).
@@ 5 Mental model improvement in surprising real-world situations
Figure 2: Deployment setup. A safety driver, a support engineer, and a researcher were present. The safety driver drove the AV manually between road tests, engaged self-driving mode at the start of each test, monitored AV performance during the test, and took over in case of unsafe driving. The support engineer deployed CW-Net and set scenario destinations. The researcher directed testing. The dashboard included a map with overlaid object detections ($s$) from the perception module and the output trajectory ($\hat{\tau}$). Explanations $\mathbf{c}_{\hat{i}}$ from CW-Net were shown as percentages for easier interpretation [39].
@@ 5.1 Unexpected stopping for nearby vehicles
Figure 3: Results: a. The CLOSE concept activated when the car got stuck next to parked vehicles. The driver initially thought the pickup/drop-off area was the cause, but the explanation suggested that it was the nearby vehicles. When the driver engaged self-driving away from the park vehicles, activation of the CLOSE concept decreased and the AV started moving again, counter to driver’s initial mental model and consistent with the explanations. Across tests, CLOSE correlated with speed and the intercept accurately predicted this event. b. The ASV concept activated when the car stopped next to a traffic cone. The driver initially thought the cone was the cause, but the explanation suggested that the AV was hallucinating a stopped vehicle. When we removed the cone in a counterfactual test, the same phantom braking and concept activation occurred. Across tests, ASV correlated with reductions in speed when spiking above 0.5 probability. c. The BIKE concept failed to activate in our first round of tests with a cyclist, but the car always stopped safely for the cyclist. After observing the explanation, the driver engaged self-driving from slower speeds in a second round of tests. Follow-up analyses revealed that the AV stopped for the cyclist due to backup safety mechanisms unrelated to CW-Net, which indicates that the driver’s increased level of caution was appropriate. a-c. (Right column) Naturalistic scenarios from public-road deployment, analogous to the scenarios observed on the private track. We used the PEDESTRIAN concept instead of the BIKE concept as it behaved similarly and pedestrians were more frequently encountered on public roads. Likewise, we used pedestrians in place of the traffic cone in our ASV tests as there were no cones encountered on the public roads. Standard error of the mean and 3-second rolling averages are shown in relevant plots.
We observed that the AV repeatedly came to a stop shortly before a pedestrian pickup/drop-off zone (Figure 3a). The driver’s intuition was that the car stopped because of the pickup/drop-off zone, but the explanations indicated that the planner stopped because it detected that it was “Close to another vehicle” (the CLOSE concept). To test this hypothesis, the driver manually moved the car farther from the parked cars. At this point, the probability of CLOSE decreased and the AV began moving again, thus supporting the alternative hypothesis. A full timeline of events is detailed in Figure 3a. We fitted the intercept of the CLOSE probability against the speed of the AV globally and found it accurately predicts stopping and starting for this event.